OpenStudy (anonymous):

find three arithmetic means between 7 and 27 a. 5, 10, 5 b. 12, 17, 22 c. 12, 19, 26 d. 15, 19, 23

OpenStudy (the_fizicx99):

Ok, I tested the options to see which made most sense and if you subtract 12-7 it will give you 5, So the difference is 5, by adding 5 + 7 = 12 12 + 5 = 17 17 + 5 = 22 22 + 5 = 27 Get it, now you can plug in it into this formula to find any term you want. f(n) = f(1) + d(n - 1) ---------------------------------------------------------------- f(6) = f(1) + 5(5- 1) f(6) = 7 + 5(4) f(6) = 7 + 20 = 27 See how that works?

OpenStudy (the_fizicx99):

That's a typo my bad, it should be f(5) = f(1) + 5(5-1)

OpenStudy (anonymous):

thank you!

OpenStudy (the_fizicx99):

np